OpenStudy (anonymous):

answer the following and show your work (5x-1)/2 + (-3x-1)/4 =9/2

OpenStudy (campbell_st):

1st multiply every term by 4 to eliminate the fractions 2(5x -1) + (-3x -1) = 2*9 10x - 2 -3x -1 = 18 collect like terms 7x - 3 = 18 add 3 to both sides 7x = 21 divide by 7 x = 3 hope it helps